# Zero-shot denoising via neural compression: Theoretical and algorithmic framework

### v1

@@ 4 Zero-shot compression-based denoiser
the first term is the noise variance, and the second is the true denoising error. While ${\bm{z}}$ and $\hat{\bm{x}}$ are not fully independent, they are intuitively weakly correlated in successful denoising regimes, where the estimate $\hat{\bm{x}}$ depends only indirectly on the noise. Thus, the cross term is expected to be small: $\frac{1}{n}\mathbb{E}\left[{\bm{z}}^{\top}({\bm{x}}-\hat{\bm{x}})\right]\approx 0.$ This approximation suggests that $\frac{1}{n}\|{\bm{y}}-\hat{\bm{x}}\|_{2}^{2}$ should be close to $\sigma_{z}^{2}$ when $\hat{\bm{x}}$ is a high-quality estimate. Based on this insight, we propose a simple and effective heuristic for choosing $\lambda$: select the value that makes $\frac{1}{n}\|{\bm{y}}-\hat{\bm{x}}\|_{2}^{2}$ closest to the known noise variance $\sigma_{z}^{2}$. This procedure can be implemented efficiently via a tree-based search strategy, as described in Algorithm 1. To apply Algorithm 1, one needs an estimate of the noise power $\sigma_{z}^{2}$. This is a well-studied problem and there exist robust algorithms for estimating the power of noise [2, 50]. For example, in [2], it is shown that the noise power can be estimated from the median of the absolute differences of wavelet coefficients. Consequently, the performance of ZS-NCD is relatively robust to the choice of $\lambda$. For instance, on the Nuclei dataset (Figure 3), ZS-NCD outperforms the state-of-the-art zero-shot learning-based denoiser, ZS-Noise2Noise, across a wide range of $\lambda$ values, for both $\sigma_{z}=10$ and $\sigma_{z}=20$. A similar approach can be used in the case of Poisson noise as well. (See Section A.2)
@@ 5 Experiments
In this section, we evaluate the denoising performance of ZS-NCD on both synthetic and real-world noise, across natural and microscopy images. We compare against representative zero-shot denoisers, including both traditional and learning-based methods. All baselines are dataset-free, i.e., they operate solely on the noisy image to be denoised. For non-learning methods, we include JPEG-2K and BM3D. Although rarely used as a denoising baseline, JPEG-2K provides a useful point of comparison from the perspective of compression-based denoising, as it represents a fixed, pre-defined compression code. For learning-based methods, we evaluate Deep Image Prior (DIP) [21], Deep Decoder (DD) [22], Zero-Shot Noise2Self (ZS-N2S) [12], Self2Self (S2S) [23], and Zero-Shot Noise2Noise (ZS-N2N) [24].
@@
We consider two synthetic noise models, AWGN $\mathcal{N}(0,\sigma^{2}_{z})$, where $\sigma_{z}$ is the standard deviation of the Gaussian distribution, and Poisson noise defined as $\mathrm{Poisson}(\alpha x)$, where $\alpha$ is the scale factor. Note that Poisson noise is signal-dependent noise with $\mathbb{E}\left[{\bm{y}}\right]=\alpha{\bm{x}}$. To re-scale the noisy image to the range of clean image, we followed the literature by assuming that the the scale $\alpha$ is known, and normalize the noisy image as ${\bm{y}}/\alpha$ in the experiments of this section. We evaluate on grayscale Set11 [25], RGB Set13 [51] (center-cropped to $192\times 192$) and Kodak24 [52] datasets. Table 1 presents the denoising performance of various methods. BM3D achieves the strongest results on grayscale images, though it relies on accurate knowledge of the noise power parameter. Existing learning-based zero-shot denoisers, in contrast, often exhibit inconsistent performance across noise levels and image resolutions. For example, ZS-N2S and Self2Self degrade on high-resolution images, likely due to the limitations of training with masked pixels. ZS-N2N performs well on high-resolution images from Kodak24 but suffers on lower-resolution images in Set13 ($192\times 192$), as it is trained to map between two downscaled versions of the same noisy image. In comparison, ZS-NCD maintains robust performance across different noise levels and image sizes. The more realistic case of not having access to the noise parameter $\alpha$ is discussed in Appendix A.2. In both noise regimes, we use MSE as the loss function. However, for Poisson noise, minimizing the negative log-likelihood is also a natural choice. We defer the results using this loss to Appendix A.2.
@@
To evaluate performance in low-data and domain-shift settings, we test ZS-NCD on Mouse Nuclei fluorescence microscopy images [53], which differ significantly from natural images in structure and texture. We also assess real-world denoising using the PolyU dataset [54], which contains high-resolution images captured by Canon, Nikon, and Sony cameras. Ground-truth images are obtained by averaging multiple captures, while the noisy inputs are single-shot acquisitions. Results are shown in Table 2. ZS-NCD consistently outperforms other learning-based zero-shot denoisers, demonstrating robustness to unknown noise models and non-natural image distributions.
Table 3: AWGN denoising on Set11.
@@
Most learning-based zero-shot methods are prone to overfitting due to the lack of clean targets and the use of overparameterized networks.
In contrast, ZS-NCD, grounded in compression-based denoising theory, overcomes this issue given the entropy constraint. To further highlight this key aspect of ZS-NCD, we replace the convolutional encoder-decoder ($\approx 0.4$M params) with a fully connected MLP ($\approx 2.3$M params) and observe that, instead of degradation, the performance improves using the same $\lambda$ (see Table 3).
@@
As described in Section 4 and illustrated in Fig. 1, ZS-NCD denoises each pixel by aggregating outputs from overlapping patches, where each patch is first compressed and then decompressed using a learned neural compression model. Intuitively, one might expect the most accurate reconstruction for a given pixel to come from the patch in which it lies at the center, as this location benefits from the largest available spatial context, which has been observed in [55].
@@
To investigate this, we conducted an ablation in which, instead of averaging, each pixel $(i,j)$ is reconstructed solely from one of the $k\times k$ patches in which it appears, using a fixed location in the patch (e.g., top-left, center, etc.). The results are shown in Fig. 4, where each heatmap entry reports the PSNR
Figure 4: Denoising AWGN ($\sigma_{z}=25$) on image Parrot using ZS-NCD, where only a single pixel from each overlapping patch (stride 1) is retained after compression. Each heatmap value indicates the PSNR achieved when denoising is based solely on the pixel at that specific location within each patch.
obtained by using only that specific location in the patch for reconstruction. As expected, performance is best when the pixel is centrally located, and degrades as it moves toward the patch boundaries.
@@
6 Proof of Theorems
6.1 Auxiliary lemmas
@@
Consider independent Poisson random variables $Y_{1},\ldots,Y_{n}$, where $Y_{i}\sim\mathrm{Poisson}(\alpha_{i})$. Consider $w_{1},\ldots,w_{n}\in\operatorname{\mathbb{R}}$. Let $\sigma_{n}^{2}=\sum_{i=1}^{n}w_{i}^{2}\alpha_{i}$ and $w_{M}\triangleq\max_{i\in\{1,\ldots,n\}}|w_{i}|$. Then, for any $t\in[0,\frac{2\sigma_{n}^{2}}{3w_{M}}]$,
($\mathbb{P}\left\lparen\sum_{i=1}^{n}w_{i}Y_{i}\geq\sum_{i=1}^{n}w_{i}\alpha_{i}+t\right\rparen\leq\exp(-{t^{2}\over 2(\sigma_{n}^{2}+w_{M}t/3)}).$ ) . (10)
@@
($\mathbb{P}\left\lparen\sum_{i=1}^{n}w_{i}Y_{i}\leq\sum_{i=1}^{n}w_{i}\alpha_{i}-t\right\rparen\leq\exp(-\frac{t^{2}}{2(\sigma_{n}^{2}+w_{M}t/3)}).$ ) . (11)
6.1.1 Proof of Lemma 1
@@
$\log(1+u)=u+f^{\prime\prime}(\alpha){u^{2}\over 2},$ (13)
@@ 6.1.2 Proof of Lemma 2
($\mathbb{P}\left\lparen\sum_{i=1}^{n}w_{i}Y_{i}\geq\sum_{i=1}^{n}w_{i}\alpha_{i}+t\right\rparen$ ) $\leq\frac{\prod_{i=1}^{n}\mathbb{E}\left[\exp(sw_{i}(Y_{i}-\alpha_{i}))\right]}{\exp(st)}$
@@
($\mathbb{P}\left\lparen\sum_{i=1}^{n}w_{i}Y_{i}\leq\mu_{n}-t\right\rparen$ ) $\leq\exp(\sum_{i=1}^{n}\alpha_{i}(e^{-sw_{i}}-1+sw_{i})-st).$ (18)
@@
($\mathbb{P}\left\lparen\sum_{i=1}^{n}w_{i}Y_{i}\leq\mu_{n}-t\right\rparen$ ) $\leq\exp\left(\frac{s^{2}\sigma_{n}^{2}}{4/3}-st\right).$ (20)
@@ 6.2 Proof of Theorem 1
where the last line follows because $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$, for all $a,b>0$. To finish the proof we need to bound ($\mathbb{P}\left\lparen(\mathcal{E}_{1}\cap\mathcal{E}_{2})^{c}\right\rparen$ ) and set parameters $t_{1}$ and $t_{2}$.
@@
($\mathbb{P}\left\lparen\left\lvert\sum^{n}_{i=1}z_{i}\frac{e_{i}^{(c)}}{\lVert{\bf e}^{(c)}\rVert}\right\rvert\geq t\right\rparen\leq 2\exp\left\lparen-\frac{t^{2}}{2\sigma_{z}^{2}}\right\rparen.$ ) . (35)
@@
($\mathbb{P}\left\lparen\mathcal{E}_{1}^{c}\cup\mathcal{E}_{2}^{c}\right\rparen\leq 2^{-\eta R+2}.$ (38)
@@ 6.3 Proof of Theorem 2
Recall that $\hat{\bm{x}}=\arg\min_{{\bf c}\in\mathcal{C}}\mathcal{L}({\bf c};{\bm{y}})$. Let

### v2

@@ 4 Zero-shot compression-based denoiser
the first term is the noise variance, and the second is the true denoising error. While $\bm{z}$} and $\hat{\bm{x}}$ are not fully independent, they are intuitively weakly correlated in successful denoising regimes, where the estimate $\hat{\bm{x}}$ depends only indirectly on the noise. Thus, the cross term is expected to be small: $\frac{1}{n}\mathbb{E}\left[{\bm{z}}^{\top}({\bm{x}}-\hat{\bm{x}})\right]\approx 0.$ This approximation suggests that, ideally, under low-noise regimes, $\frac{1}{n}\|{\bm{y}}-\hat{\bm{x}}\|_{2}^{2}$ is expected to be close to $\sigma_{z}^{2}$. Based on this insight, we propose a simple and effective heuristic for choosing $\lambda$: select $\lambda$ such that $\frac{1}{n}\|{\bm{y}}-\hat{\bm{x}}\|_{2}^{2}$ is closest to the known noise variance $\sigma_{z}^{2}$. This procedure can be implemented efficiently via a tree-based search strategy, as described in Algorithm 1. To apply Algorithm 1, one needs an estimate of the noise power $\sigma_{z}^{2}$. This is a well-studied problem and there exist robust algorithms for estimating the variance of noise [donoho1994ideal, chen2015efficient]. For example, in [donoho1994ideal], it is shown that the noise power can be estimated from the median of the absolute differences of wavelet coefficients.
Finally, we observe that the performance of ZS-NCD is relatively robust to the choice of $\lambda$. For instance, on the Nuclei dataset (Figure 3), ZS-NCD outperforms the state-of-the-art zero-shot learning-based denoiser, ZS-Noise2Noise, across a wide range of $\lambda$ values, for both $\sigma_{z}=10$ and $\sigma_{z}=20$. A similar approach can also be applied to the case of Poisson noise, as well.
Case II: Poisson noise In the case of Poisson noise, in addition to estimating $\lambda$, we need to estimate $\alpha$, which is used to normalize the measurements, in both the MSE-based and the MLE-based methods. Note that in the case of Poisson noise, $\mathbb{E}[y_{i}]=\alpha x_{i}$, and therefore, with high probability, $\frac{1}{n}\sum^{n}_{i=1}y_{i}\approx\alpha\frac{1}{n}\sum^{n}_{i=1}x_{i}$. Using this observation, and assuming that $\frac{1}{n}\sum^{n}_{i=1}x_{i}\approx 0.5$, we estimate $\alpha$ as $\hat{\alpha}=\frac{2}{n}\sum^{n}_{i=1}y_{i}$. We then use the estimated noise level $\hat{\alpha}$ to normalize both MSE and MLE based optimization for denoising Poisson noise. See Table 1 for the result of this noise parameter estimation on a sample image. Having an estimate of noise parameter $\alpha$, we then follow a similar procedure we used in the case of AWGN to set the parameter $\lambda$. Specifically, we write the MSE between normalized $\bm{y}$} and the denoised image $\hat{\bm{x}}$ as
[…]
Again, assuming that we are in a low-noise regime, i.e., the second and the third terms in (11) are close to zero, and using ${1\over n}\sum_{i=1}^{n}x_{i}\approx 0.5$ approximation, it follows that ${1\over n}\mathbb{E}\left[\|{\bm{y}}/\alpha-\hat{\bm{x}}\|_{2}^{2}\right]\approx\frac{1}{2\alpha}$. This implies that, in the case of Poisson noise, we can still use Algorithm 1 to find $\lambda$, after updating $\|\hat{\bm{x}}^{(k)}-\bm{y}\|^{2}>n\sigma_{z}^{2}$ to $\|\hat{\bm{x}}^{(k)}-\bm{y}\|^{2}>\frac{1}{2\alpha}$. We empirically observe that MSE(${\bm{y}}/\alpha$$\hat{\bm{x}}$) in training our networks is close to $\frac{1}{2\alpha}$ as reported in Table 1, which indicates that $\frac{1}{2\alpha}$ is a good approximation of the MSE that can be used as a threshold for selecting $\lambda$. When $\alpha$ is not known, we obtain the estimate $\hat{\alpha}$, and use $\frac{1}{2\hat{\alpha}}$ as a valid threshold to set $\lambda$.
true $\alpha$ estimated $\hat{\alpha}$ empirical MSE(${\bm{y}}/\alpha$$\hat{\bm{x}}$) [dB] $1/(2\alpha)$ $1/(2\hat{\alpha}) [dB$]
25 23.02 17.12 16.98 16.63
50 46.05 19.66 20.00 19.64
@@ 5 Experiments
In this section, we evaluate the denoising performance of ZS-NCD on both synthetic and real-world noise, across natural and microscopy images. We compare against representative zero-shot denoisers, including both traditional and learning-based methods. All baselines are dataset-free, i.e., they operate solely on the noisy image to be denoised. For non-learning methods, we include JPEG-2K and BM3D. Although rarely used as a denoising baseline, JPEG-2K provides a useful point of comparison from the perspective of compression-based denoising, as it represents a fixed, pre-defined compression code. For learning-based methods, we evaluate Deep Image Prior (DIP) [dmitry2020dip], Deep Decoder (DD) [heckel2018deep], Zero-Shot Noise2Self (ZS-N2S) [batson2019noise2self], Self2Self (S2S) [quan2020self2self], and Zero-Shot Noise2Noise (ZS-N2N) [mansour2023zsn2n].
@@
We consider two synthetic noise models, AWGN $\mathcal{N}(0,\sigma^{2}_{z})$, where $\sigma_{z}$ is the standard deviation of the Gaussian distribution, and Poisson noise defined as $\mathrm{Poisson}(\alpha x)$, where $\alpha$ is the scale factor. Note that Poisson noise is signal-dependent noise with $\mathbb{E}\left[{\bm{y}}\right]=\alpha{\bm{x}}$. To re-scale the noisy image to the range of clean image, we followed the literature by assuming that the the scale $\alpha$ is known, and normalize the noisy image as ${\bm{y}}/\alpha$ in the experiments of this section. We evaluate on grayscale Set11 [zhang2017beyond], RGB Set13 [zeyde2010single] (center-cropped to $192\times 192$) and Kodak24 [kodak1993kodak] datasets. Table 2 presents the denoising performance of various methods. BM3D achieves the strongest results on grayscale images, though it relies on accurate knowledge of the noise power parameter. Existing learning-based zero-shot denoisers, in contrast, often exhibit inconsistent performance across noise levels and image resolutions. For example, ZS-N2S and Self2Self degrade on high-resolution images, likely due to the limitations of training with masked pixels. ZS-N2N performs well on high-resolution images from Kodak24 but suffers on lower-resolution images in Set13 ($192\times 192$), as it is trained to map between two downscaled versions of the same noisy image. In comparison, ZS-NCD maintains robust performance across different noise levels and image sizes. The more realistic case of not having access to the noise parameter $\alpha$ was discussed in Section 4. In both noise regimes, we use MSE as the loss function. However, for Poisson noise, minimizing the negative log-likelihood is also a natural choice. We defer the results using this loss to Appendix B.3.
@@
To evaluate performance in low-data and domain-shift settings, we test ZS-NCD on Mouse Nuclei fluorescence microscopy images [buchholz2020denoiseg], which differ significantly from natural images in structure and texture. We also assess real-world denoising using the PolyU dataset [xu2018real], which contains high-resolution images captured by Canon, Nikon, and Sony cameras. Ground-truth images are obtained by averaging multiple captures, while the noisy inputs are single-shot acquisitions. Results are shown in Table 3. ZS-NCD consistently outperforms other learning-based zero-shot denoisers, demonstrating robustness to unknown noise models and non-natural image distributions.
Table 4: Comparison of AWGN denoising of Conv and MLP based ZS-NCD on Set11.
@@
Most learning-based zero-shot methods are prone to overfitting due to the lack of clean targets and the use of overparameterized networks. In contrast, ZS-NCD, grounded in compression-based denoising theory, overcomes this issue given the entropy constraint. To further highlight this key aspect of ZS-NCD, we replace the convolutional encoder-decoder ($\approx 0.4$M params) with a fully connected MLP ($\approx 2.3$M params) and observe that, instead of degradation, the performance improves using the same $\lambda$ (see Table 4).
@@
As described in Section 4 and illustrated in Fig. 1, ZS-NCD denoises each pixel by aggregating outputs from overlapping patches, where each patch is first compressed and then decompressed using a learned neural compression model. Intuitively, one might expect the most accurate reconstruction for a given pixel to come from the patch in which it lies at the center, as this location benefits from the largest available spatial context, which has been observed in [he2021checkerboard].
@@
To investigate this, we conducted an ablation in which, instead of averaging, each pixel $(i,j)$ is reconstructed solely from one of the $k\times k$ patches in which it appears, using a fixed location in the patch (e.g., top-left, center, etc.). The results are shown in Fig. 4, where each heatmap entry reports the PSNR obtained by using only that specific location in the patch for reconstruction. As expected, performance is best when the pixel is centrally located, and degrades as it moves toward the patch boundaries.
@@
Figure 4: Denoising Parrot using ZS-NCD, where only a single pixel from each overlapping patch (stride 1) is retained after compression. (AWGN, $\sigma_{z}=25$.) Each heatmap value indicates the PSNR achieved when denoising is based solely on the pixel at that specific location within each patch.
6 Conclusions
We have studied maximum likelihood compression-based denoising, and provided theoretical characterization of its performance under both AWGN and Poisson noise. Furthermore, we introduced ZS-NCD, a new zero-shot neural-compression-based denoising and demonstrated that it achieves state-of-the-art performance among zero-shot methods, in both AWGN and Poisson denoising.
The presented theoretical results are derived by assuming a fixed (e.g., pre-trained/defined) compression code. Extending these results to the case of zero-shot learned compression codes is an interesting direction for future research.
Acknowledgment
A.Z., X.C., S.J. were supported by NSF CCF-2237538.
Appendix A Proofs
A.1 Auxiliary lemmas
@@
Consider independent Poisson random variables $Y_{1},\ldots,Y_{n}$, where $Y_{i}\sim\mathrm{Poisson}(\alpha_{i})$. Consider $w_{1},\ldots,w_{n}\in\mathbb{R}$. Let $\sigma_{n}^{2}=\sum_{i=1}^{n}w_{i}^{2}\alpha_{i}$ and $w_{M}\triangleq\max_{i\in\{1,\ldots,n\}}|w_{i}|$. Then, for any $t\in[0,\frac{3\sigma_{n}^{2}}{2w_{M}}]$,
($).\displaystyle\mathbb{P}\left\lparen\sum_{i=1}^{n}w_{i}Y_{i}\geq\sum_{i=1}^{n}w_{i}\alpha_{i}+t\right\rparen\leq\exp(-{t^{2}\over 2(\sigma_{n}^{2}+w_{M}t/3)}$). (12)
@@
($).\displaystyle\mathbb{P}\left\lparen\sum_{i=1}^{n}w_{i}Y_{i}\leq\sum_{i=1}^{n}w_{i}\alpha_{i}-t\right\rparen\leq\exp(-\frac{t^{2}}{2(\sigma_{n}^{2}+w_{M}t/3)}$). (13)
A.1.1 Proof of Lemma 1
@@
$\log(1+u)=u-f^{\prime\prime}(\alpha){u^{2}\over 2},$ (15)
@@ A.1.2 Proof of Lemma 2
[…]
@@
[…]
@@
[…]
@@ A.2 Proof of Theorem 1
[…]
@@
[…]
@@
[…]
@@ A.3 Proof of Corollary 1
We start by designing a lossy compression code for the set of signals in $\mathcal{Q}_{n}$, defined as
$\mathcal{Q}_{n}=\left\{{\bm{x}}\in\operatorname{\mathbb{R}}^{n}:\lVert{\bm{x}}\rVert_{0}\leq k,\lVert{\bm{x}}\rVert_{\infty}\leq 1\right\}.$ (41)
For a $k$-sparse ${\bm{x}}\in\mathcal{Q}_{n}$, let ${\bm{x}}^{(k)}\in\operatorname{\mathbb{R}}^{k}$ denote the $k$-dimensional vector derived from the non-zero coordinates of $\bm{x}$}. We define a lossy compression code $(f,g)$ that achieves distortion $\delta$. Specifically, given a $k$-sparse ${\bm{x}}\in\mathcal{Q}_{n}$, the encoder operates as follows:
  1. 1.
    Encode the number of non-zero entries and their locations. This requires at most
    $\log_{2}k+\log_{2}{n\choose k}\leq\log_{2}k+k\log_{2}({n{\rm e}\over k})$
    bits.
  2. 2.
    Quantize the values of the non-zero coordinates such that overall distortion $\delta$ is achieved. To achieve this goal it quantizes each non-zero coordinate of $\bm{x}$} into $b$ bits. Let $x_{i}]_{b}$ denote the $b$-bit quantized version of $x_{i}$. Then, $|x_{i}-[x_{i}]_{b}|\leq 2^{-b}$. Therefore, the overall $\ell_{2}$ distortion can be bounded as
    $\|{\bm{x}}-g(f({\bm{x}}))\|_{2}^{2}\leq k2^{-2b}.$
    Choosing
    $b=\lceil{1\over 2}\log_{2}{k\over n\delta}\rceil$
    ensures that $\|{\bm{x}}-g(f({\bm{x}}))\|_{2}^{2}\leq n\delta$.
It can be observed that overall the number of bits required for describing the signals in $\mathcal{Q}_{n}$ within distortion $\delta$ can be bounded as
$R\leq{k\over 2}\log_{2}({k\over n\delta})+k\log_{2}{n\over k}+c_{k},$
where $c_{k}=\log_{2}k+k(\log_{2}{\rm e}+1)$. Using the defined lossy compression code to solve (1) and applying Theorem 1, it follows that, with a probability larger than $1-2^{-\eta R+2}$,
${1\over\sqrt{n}}\|{\bm{x}}-\hat{\bm{x}}\|_{2}$ $\leq\sqrt{\delta}+\sigma_{z}C\sqrt{{k\over 2n}\log_{2}({k\over n\delta})+{k\over n}\log_{2}({n\over k})+{c_{k}\over n}}$
$=\sqrt{\delta}+\sigma_{z}C\sqrt{{k\over 2n}\log_{2}({n\over k\delta})+{c_{k}\over n}},$
where $C=2(1+2\sqrt{\eta})\sqrt{2\ln 2}$. Let
$\delta={1\over n}.$
Then,
${1\over\sqrt{n}}\|{\bm{x}}-\hat{\bm{x}}\|_{2}$ $\leq{1\over\sqrt{n}}+\sigma_{z}C\sqrt{{k\log_{2}k\over 2n}+{k\over n}\log_{2}({n\over k})+{c_{k}\over n}}$
$={1\over\sqrt{n}}+\sigma_{z}C\sqrt{{k\over n}\log_{2}({n\over k})+\gamma_{n}},$
where
$\gamma_{n}={k\log_{2}k\over 2n}+{c_{k}\over n}.$
Finally, since $R\geq k\log_{2}({n\over k})$,
$1-2^{-\eta R+2}\geq 1-{\rm e}^{-\eta k\log({n\over k})}.$
A.4 Proof of Theorem 2
Proof.
Recall that $\arg\min_{{\bf c}\in\mathcal{C}}\mathcal{L}({\bf c};{\bm{y}})$. Let